The length of each side of a square is extended 2 inches. The area of the resulting square is 144 square inches. Find the length of the side of the original square.
Answer by jim_thompson5910(35256)

(x+2)^2 = 144

x^2+4x+4=144

x^2+4x+4-144=0

x^2+4x-140

(x+14)(x-10)=0

x+14=0 or x-10=0

x=-14 or x=10

Ignore the negative solution.

So the length of the side of the original square is 10 inches.